kittybasil:

Calculus I Question! "5.2: The Definite Integral" "Use the Midpoint Rule with the given value of n to approximate the integral. Round the answer to four decimal places."

kittybasil:

\[\large{\int^{x}_{0}e^{x}\cos{x}dx}\]\[n=4\]

Angle:

\(\huge \int^{n}_{0}\) ??

kittybasil:

Eh?

Angle:

an x or n there?

kittybasil:

OH SHOOT! That's supposed to be a \(\pi\). LOL my bad

Angle:

ohhhhh that would make a ton more sense

kittybasil:

\[\Large{\int^{\pi}_{0}e^{x}\cos{x}dx}\]\[n=4\]

Angle:

:scrunches face: I'm guessing n = number of rectangles they want us to use?

kittybasil:

I think \(n=\text{sub-intervals}\)

Angle:

yeahhh number of rectangles, basically midpoints tho >.> it's called midpoint, because the height is the middle of the rectangle 

kittybasil:

So basically like that last question, but using midpoint of the rectangle width?

Angle:

base = pi/4 height = f( the midpoints ) f(x) being = e^x * cos(x) yes :)

kittybasil:

Hey, sorry for disappearing on you. I had to move rooms since my college has wonky closing hours for certain locations... Not sure if you got my PM? @Angle

Angle:

no worries ^_^ you know how to work through the rest? or are you still confused at some part? :3

kittybasil:

I'm not sure right now... maybe get back to you later on that? ^^;

Angle:

sure thing ^_^

kittybasil:

Alright cool :D *Medal system be trolling right now. D: